Determine the [OH−] , pH, and pOH of a solution with a [H+] of 0.0029 M at 25 °C. [OH−]= M pH= pOH= Determine the [H+] , pH, and pOH of a solution with an [OH−] of 3.4×10−8 M at 25 °C. [H+]= M pH= pOH= Determine the [H+] , [OH−] , and pOH of a solution with a pH of 3.93 at 25 °C. [H+]= M [OH−]= M pOH= Determine the [H+] , [OH−] , and pH of a solution with a pOH of 4.35 at 25 °C. [H+]= M [OH−]= M pH=
Answer #1

TH+] = 0.0029 Date 11 Page at 251 pH = - wg [ft] = - loe [0.0029] pH = 2.53 poht pH = 14 POH= 14-2.5} = 11.47 [OH- ] = 10-14